OpenStudy (anonymous):

Find the critcal points and using the second derivative test determine if local max/min exists or if the test is inconclusive. x^4/2-12x^2

OpenStudy (shamil98):

\[\huge \frac{ x^4 }{ 2 } - 12x^2\] is this your problem?

OpenStudy (anonymous):

When i took both derivatives i got,\[f \prime = 2x^3-24\] and \[f \prime \prime\] = 6x^2-24\]

OpenStudy (anonymous):

^^yes

OpenStudy (shamil98):

nvm, you made a typo lol

OpenStudy (anonymous):

I factored out the first derivative to get, \[2x(x^2-12)=0\]. Then i plugged in my points , -4,-1,2,and 4 in the 2nd derivative to get, a local max at -4, -1 and local min at 2, and 4

OpenStudy (anonymous):

oh i forgot my critical points are \[x=0,\sqrt{12}\]

OpenStudy (shamil98):

yeah, mb , i was looking at the second derivative for a sec, you're right with the critical points

OpenStudy (anonymous):

so i just want to make sure that when i'm solving these, I must always factor out the first derivative not the second one correct?

OpenStudy (shamil98):

yeah, i made mistake looking at the second derivative, the critical points are found by taking the first derivative and seeing which values of x make the derivative zero

OpenStudy (anonymous):

oh ok